Sol Summary – April 25th

Crew 280 Sol Summary Report 25-04-2023

Sol: 9

Summary title: Martian Days, Universe-Sized Dreams

Author’s name: Mariona Badenas-Agustí, Commander.

Mission status: Nominal

Sol activity summary:

Last night, Mariona Badenas-Agustí (Commander & Astronomer) led an astronomy workshop to share her love for the Universe with the Hypatia I crew. This activity sparked conversations about the immensity and diversity of the cosmos, as well as about the power of science to drive progress and social change.

Following up on these themes, the crew began SOL9 with a lively breakfast discussion about personal career aspirations. Each crew member shared their dreams for the future before resuming their experimental work. With the end of the mission in sight, everyone is feeling the intensity of the final stretch.

During the morning, the Hab and the Science Dome were filled with a quiet atmosphere of silence and concentration. Experiments are currently in full swing, and as it always happens in science, results don’t always go well and challenges have become a daily habit.

For lunch, Mariona and Carla Conejo González (Executive Officer & Biologist) went creative and made a delicious meal of pasta with tomato sauce and a fresh green salad. In the afternoon, Ariadna Farrés (Scientist & HSO) and Neus Sabaté (Mission Specialist) had an EVA planned to collect more iron-rich rock batteries, which are already giving promising results. Meanwhile, the crew members who stayed in the Hab continued to work on their experiments.

Tonight, the crew is thinking about watching a documentary after enjoying another Martian dinner at the MDRS.

Look ahead plan: There will be one morning EVA, led by Mariona Badenas, Carla Conejo, Núria Jar, and Laia Ribas (see EVA Request).

Anomalies in work: None.

Weather: Sunny. Low 45F/High 67F.

Crew physical status: Nominal.

EVA: There was one EVA today to Barrainca Butte(see EVA Report).

Reports to be filed: Sol Summary, EVA request, EVA Report, Operations Report, GreenHab Report, Journalist Report, Pictures of the Day.

Support Requested: None.

Operations Report – April 25th

Crew 280 Operations Report 25-04-2023
SOL: 9
Name of person filing report: Cesca Cufi Prat (Crew Engineer)
Non-nominal systems:
Robotic Observatory
Notes on non-nominal systems: Replacement parts of Robotic observatory are scheduled to be installed during crew 261 rotation.
Spirit rover used: yes
Hours: 218.6
Beginning charge: 100%
Ending charge: 67%
Currently charging: no
Opportunity rover used: no
Hours: 121.9
Beginning charge: –
Ending charge: –
Currently charging: yes
Curiosity rover used: no
Hours: 231.0
Beginning charge: –
Ending charge: –
Currently charging: yes
Perseverance rover used: no
Hours: 262.0
Beginning charge: –
Ending charge: –
Currently charging: yes
General notes on rovers: none
Summary of Hab operations:
WATER USE: 30,6 gal (from 6PM Sol4 to 6PM Sol5)
Water (static tank): 290 gal (at 6:00PM)
Static tank pipe heater: off
Static tank heater: off
Toilet tank emptied: no
Summary of internet: used from 18:00 to 21:00 for personal communication and reports.
Summary of suits and radios:
Suits: 3 and 6.
Radios: nominal
All radio and spacesuit batteries checked before EVA: Yes. All nominal.
Summary of GreenHab operations:
WATER USE: 12 gallons
Heater: Off
Supplemental light: Off
Harvest: 395 g microgreen salad, 99 g of red onions, 40 g chives, 60 g jazz mix, 4 g microgreen, 2 g salvia, 1 g tyme, 1 g albahaca (sweet basil).
Summary of ScienceDome operations: Laia Ribas, Neus Sabaté and Cesca Cufí have been working on their experiments.
Dual split: Heat or AC, Off
Summary of RAM operations: use of solder to repair (for the second time) Laia Ribas’ Martian Bottle.
Summary of any observatory issues:
Robotic observatory: Not working. Replacement parts are scheduled to be installed during crew 261 rotation.
Summary of health and safety issues: none
Summary of other system’s issues:
Heater: stop due to wind. Restarted as per provided procedure.
Questions, concerns and requests to Mission Support:
Questions: none
Kitchen: knobs replacement – expected this week
Retrieval of trash in the rear airlock

Journalist Report – April 25th

Crew 280 Journalist Report 25-04-2023

Author: Núria Jar, Crew Journalist

The wind is a constant presence at the Mars Desert Research Station (MDRS). The sound of the canvas flapping against the corridors of the MDRS serves as a constant background melody for the crew living inside the Hab. Dust storms are common on Mars and they can last for weeks, sometimes even months, covering the surface of the Red Planet. Precisely, this is one of the most important meteorological phenomena of Mars.

Despite the high winds, two crew members from Hypatia I conducted an extravehicular activity (EVA) because it was deemed safe to do so. Whether staying inside the base or going out, preparing for an EVA is always a special moment for the crew members. They have to help each other in suiting up, checking the equipment and communication effectively before heading out. Playing music through the radios has quickly become a tradition while the crew members who are about to conduct an EVA are waiting in the airlock to depressurize. Five minutes can feel like an eternity in the cramped space of an airlock, sandwiched between two doors.

While the two crew members were outside conducting their EVA, the remaining crew stayed inside the base to continue their ongoing experiments. One such experiment is called the “Martian bottle”, which aims to recycle water as a means of promoting self-sustainability. Water is a scarce resource on Mars, and the crew is trying to generate drinkable water from the used water in the GreenHab using the bottle cap Light Pills, designed by Hypatia I crew member Helena Arias.

Throughout the day, there were delayed communications between Mars and Earth. This afternoon, the two back-up crew members of Hypatia I, Helena Arias and Anna Bach, were busy in Barcelona discussing the initial days of the mission at the MDRS. They shared their insights and impressions about the mission at a gathering event held at the Fabra Observatory. The event was a success and people in Catalonia are interested in learning more about the Hypatia I mission, but the delay between the two planets is challenging.

Meanwhile, the crew members shared their dreams for the future after their experience at the MDRS. The crew’s daily routine at the station is governed by the rhythm of Mars, yet their aspirations reach far beyond. Ad Astra, Per Aspera!

GreenHab Report – April 25th

Crew 280 Green Hab Report 25-04-2023

SOL 9 Greenhab Report

GreenHab Officer: Laia Ribas

Environmental control: Open during the day and heater during the night

Temperatures: 74.4 F

Hours of supplemental light: N/A

Daily water usage for crops: 12 gallons

Daily water usage for research and/or other purposes: 0 gallons

Water in Blue Tank: 79 gallons

Time(s) of watering for crops: 08.00, 18:00 h

Changes to crops: N/A

The general watering of plants and crops was done. The door was closed in the morning due to wind and opened during the afternoon to decrease the temperature. The temperature reached 84.5F at 14.00h.

Harvest: 395 g microgreen salad, 99 g of red onions, 40 g chives, 60g jazz mix, 4g microgreen, 2g salvia, 1g tyme, 1g albahaca (sweet basil).

Based on Mission Support’s request, today the red onions (planted in 2022) have been all harvested. The chives have been partially harvested (about one-third). The blue tub of microgreen salads has been completely harvested and replanted by using two different kinds of seeds in half of the container (Miners lettuce and Mustard Spinach). More soil from a bag in the GreenHab has been added together with a scoop of Plan Food (Miracle Gro) to fertilize. See pictures of the current replant.

Support/supplies needed.
The black pen marker belonging to the Green Hab doesn’t write properly.

EVA Report – April 25th

Crew 280 EVA Report 25-04-2023

EVA #9

Author: Ariadna Farres (HSO & Scientist)

The EVA had the following goal:

Collect some iron-containing rock samples to evaluate its use for Martian batteries by Neus Sabate.

Start time: 2:35 PM

End time: 5:10 PM

Narrative: Before starting the EVA, we received a call from mission support advising us of the possibility of rain in the next few hours. During the whole EVA the astronauts checked the weather, which ended up being windy but with no signs of rain and the mission could be completed.

After performing 5 minutes of depressurization in the front airlock, the EVA team took Spirit rovers and headed White Rock Canyon and Barrainca Butte area to collect iron-containing rocks as suggested by Mission Support Dr. Shannon Rupert.

The EVA team drove very slowly along Cow Dung Road to decrease the discharging rate of the battery rovers. Once they reached the selected area they observed a larger amount of rocks around Barainca Butte. They spent about an hour exploring the area and collecting rocks. At 4:30PM the crew decided to head back to the base, which they reached at 5:10PM. Before entering the Hab, they followed the 5-minutes pressurization protocol.

Destination: Barainca Butte

Coordinates (use UTM WSG 84): 520000 E, 4248000 N

EVA participants: 2: Ariadna Farrés (HSO and EVA Leader) and Neus Sabate (Mission Specialist).

Road(s) and routes per MDRS Map: Drive south on Cow Dung Road 0110 and stop after the junction with Watney Road.

Mode of travel: Drive south on Cow Dung Road 0110 and stop after the junction with Watney Road. Walked around Barainca Butte. Then walk around no farther than 1500m from the rovers.

Copyright © The Mars Society. All rights reserved. | Main Site